log x - log(x+1)= 4
and
ln 3 + ln x= 15

Answer by stanbon(48535)

log x - log(x+1)= 4
log[x/(x+1)]=4
x/(x+1)=10^4
x=10000x+10000
9,999x=-10000
x=10000/9999
=============
Cheers,
Stan H.